# A CLIP-Powered Framework for Robust and Generalizable Data Selection

### v1

@@ Abstract
Large-scale datasets have been pivotal to the advancements of deep learning models in recent years, but training on such large datasets invariably incurs substantial storage and computational overhead. Meanwhile, real-world datasets often contain redundant and noisy data, imposing a negative impact on training efficiency and model performance. Data selection has shown promise in identifying the most representative samples from the entire dataset, which aims to minimize the performance gap with reduced training costs. Existing works typically rely on single-modality information to assign importance scores for individual samples, which may lead to inaccurate assessments, especially when dealing with noisy or corrupted samples. To address this limitation, we propose a novel CLIP-powered data selection framework that leverages multimodal information for more robust and generalizable sample selection. Specifically, our framework consists of three key modules—dataset adaptation, sample scoring, and selection optimization—that together harness extensive pre-trained multimodal knowledge to comprehensively assess sample influence and optimize the selection results through multi-objective optimization. Extensive experiments demonstrate that our approach consistently outperforms existing state-of-the-art baselines on various benchmark datasets. Notably, our method effectively removes noisy or damaged samples from the dataset, enabling it to achieve even higher performance with less data. This indicates that it is not only a way to accelerate training but can also improve overall data quality. The implementation will be made publicly available soon.
@@ 1 Introduction
In this paper, we propose a CLIP-powered data selection framework that employs multimodal information for more robust and generalizable data selection, where the category text serves as strong complements to the image modality and promotes the overall performance. The framework consists of three modules, namely dataset adaptation, sampling scoring, and selection optimization module. First, the dataset adaptation module integrates image and text adapters to facilitate the transfer of pretraining knowledge to the target data. Subsequently, the sample scoring module calculates the Semantic Alignment Score (SAS) and Sample Diversity Score (SDS) based on the adapted multimodal features, which measure the image-text alignment and the variability of local patterns. Using these two scores together can select semantically representative samples while maintaining their inherent diversity. Further, in order to address the group effect, we introduce a selection optimization module to identify the ideal subsets w.r.t. the expected selection ratio through a multi-objective optimization strategy. By leveraging the multi-modal information and carefully designed supervision signals, our framework enables the selection of high-quality samples in a flexible and efficient manner.
Comprehensive evaluation across various benchmark datasets demonstrates that our approach effectively improves the performance of selected datasets, especially on large-scale datasets such as ImageNet-1k Deng et al. (2009). Moreover, the selected datasets exhibit superior cross-architecture generalization across ResNet-18/50 He et al. (2016), ViT Dosovitskiy et al. (2020), VGG-16 Simonyan & Zisserman (2014), etc. Notably, since most existing methods are not robust to more complex and realistic scenarios, we further demonstrate the strong robustness of our approach in these more challenging scenes. For instance, our proposed method can achieve an 8.13% improvement in accuracy on CIFAR-100 and a 4.41% improvement on Tiny-ImageNet compared to the leading baselines. Meanwhile, superior performance is achieved with very high efficiency compared to other baselines.
@@ 2 Related Work
Date-efficient deep learning generally incorporates dynamic data pruning Qin et al. (2024); Raju et al. (2021), data distillation Lei & Tao (2023); Du et al. (2023); Zhang et al. (2023); Cazenavette et al. (2022), data condensation Liu et al. (2023); Yang et al. (2023a); Zhao & Bilen (2021), and static data selection Xia et al. (2023); Tan et al. (2024); Yang et al. (2023b). Following the static data selection, we propose a method capable of identifying representative and diverse samples across various selection ratios. Data selection, or static dataset pruning, aims to identify and select the most representative samples from training datasets. Training on these selected samples can achieve comparable performance to that obtained with the full dataset while reducing training and storage costs. Current data selection methods can be broadly divided into three categories: importance criteria Paul et al. (2021); Tan et al. (2024), dataset distribution-based methods Xia et al. (2023); Zheng et al. (2023), and optimization-based methods Nohyun et al. (2023); Yang et al. (2023c).
@@
Dataset distribution-based methods select samples by analyzing the geometric distribution of datasets. For instance, Herding Welling (2009) determines the sample importance according to samples’ distance to their corresponding class centers. The work Ramalingam et al. (2023) applies greedy k-center to select the coreset with good data coverage. D2 Maharana et al. (2023) calculates and updates the difficulty scores of each sample by incorporating the difficulty of its neighboring examples. Moderate-DS Xia et al. (2023) chooses samples with closer distances to the median score, while CCS Zheng et al. (2023) balances the data distribution and the sample importance in selection.
Optimization-based methods have been proposed to select samples through various optimization techniques Tan et al. (2024), such as gradient matching Mirzasoleiman et al. (2020b); Killamsetty et al. (2021a), scalable self-supervised pruning metric Sorscher et al. (2022), influence function Yang et al. (2023b); Pooladzandi et al. (2022), bi-level optimization Killamsetty et al. (2021b), facility location function Mirzasoleiman et al. (2020a); Yang et al. (2023d), and submodularity Nohyun et al. (2023); Kothawade et al. (2022); Wei et al. (2015). In contrast to these methods that only rely on image information, we leverage multimodal messages for data selection, which incorporates the semantic alignment between image data and corresponding category information. This contributes to comprehensive assessments of sample effectiveness, particularly in complex scenarios where samples may be corrupted or wrongly labeled.
@@ 3.2 Sample Scoring
For selected datasets, the reduced data volume may limit the diversity of the selected data, which is important for training datasets Yang et al. (2024). To address this, we introduce another diversity perspective to comprehensively assess the effect of training data. The Sample Diversity Score (SDS) is defined as the average distance between each sample and its $k$ neighbor samples of the same class:
@@ 3.3 Selection Optimization
where $\mathbb{I}$ is an indicator function, and $s_{r}$ denotes the expected selection ratio. The loss term guides the parameter $\bm{d}$ toward near-binary values, ensuring the count of ones aligns with the expected sample size. Since the selection is guided by adaptive optimization, the final selection ratio may deviate slightly from the target. To minimize this deviation, we constrain $\mathcal{L}_{s}$ with a threshold $\theta$, which in our work is set to $5\times 10^{-4}$, ensuring the actual selection ratio differs by less than ±$\pm 0.05\%$ from the expected value. Finally, the overall loss function is formulated as:
@@
where $\alpha$ and $\beta$ are coefficients that adjust for numerical differences among the loss terms and can be set conveniently.
@@
The proposed method comprises three main components. 1) The dataset adaptation involves fine-tuning the image and text adapters. Since the adapters consist of simple linear layers, the number of parameters is small, and both the forward passes and backward passes are computationally efficient. 2) In the sample scoring process, the complexity of calculating the SAS and SDS is $O(N)$ and $O(K*d)$, respectively, where $K$ is the number of categories and $d$ is the feature dimension, typically 512. The complexity of the KNN algorithm is $O(N_{k}*d)$, where $N_{k}$ is the number of samples per class. Given that $K$ and $d$ are constants and $N_{k}$ is usually much smaller than $N$, the overall complexity of this process is approximately $O(N)$. 3) The selection optimization of $\bm{d}$ does not involve deep models and is a numerical optimization process. The complexity is proportional to the number of parameters, i.e., $O(|w|)=O(N)$.
4 Experiment
4.1 Experimental Setup
Datasets and network architectures. Consistent with previous works Tan et al. (2024); Xia et al. (2023); Zheng et al. (2023), we evaluate the effectiveness of our proposed method on various popularly used benchmark datasets, including CIFAR-10/100 Krizhevsky et al. (2009), Tiny-ImageNet Chrabaszcz et al. (2017), and ImageNet-1k Deng et al. (2009). To evaluate the generalization performance of our selected datasets, we study the effectiveness of our proposed method on a wide range of network architectures, including ResNet-18/50 He et al. (2016), Vision Transformer (ViT) Dosovitskiy et al. (2020), Swin-Transformer Liu et al. (2021), VGG-16 Simonyan & Zisserman (2014), and DenseNet-121 Huang et al. (2017).
Parameter settings. The parameters in our proposed method can be easily set. The coefficient $\alpha$ is proportional to the expected selection rate $s_{r}$, balancing the importance of dataset diversity. For instance, $\alpha$ can be set equivalent to $s_{r}$. The coefficient $\beta$ is set to 2 on all other datasets to adjust the numerical differences among loss items.
@@ 4.2 Comparison with the State-of-the-arts
Table 2: Experimental results (accuracy, %) on CIFAR-100 and Tiny-ImageNet with noisy labels. 20% of labels are disturbed. We report mean $\pm$ standard deviation.
Method / Selection Ratio CIFAR-100 (label noise) Tiny-ImageNet (label noise)
20% 30% 20% 30%
Random 34.47$\pm$0.64 43.26$\pm$1.21 17.78$\pm$0.44 23.88$\pm$0.42
MoSo 31.01$\pm$0.67 43.73$\pm$0.14 21.55$\pm$0.37 27.80$\pm$0.16
Moderate-DS 40.25$\pm$0.12 48.53$\pm$1.60 19.64$\pm$0.40 24.96$\pm$0.30
Glister 28.51$\pm$1.46 43.16$\pm$1.31 21.61$\pm$0.19 25.45$\pm$0.23
Herding 42.29$\pm$1.75 50.52$\pm$3.38 18.98$\pm$0.44 24.23$\pm$0.29
Forgetting 36.53$\pm$1.11 45.78$\pm$1.04 13.20$\pm$0.38 21.79$\pm$0.43
GraNd 31.72$\pm$0.67 42.80$\pm$0.30 18.28$\pm$0.32 23.72$\pm$0.18
EL2N 29.82$\pm$1.19 33.62$\pm$2.35 13.93$\pm$0.69 18.57$\pm$0.31
Self-sup. prototypes 31.08$\pm$0.78 41.87$\pm$0.63 15.10$\pm$0.73 21.01$\pm$0.36
CG-Score 6.82$\pm$1.60 20.07$\pm$0.45 8.35$\pm$0.65 15.31$\pm$0.90
Ours 45.63$\pm$0.34 58.65$\pm$0.46 25.98$\pm$0.19 32.21$\pm$0.24
Table 3: Numerical analysis on the noisy data ( proportion, %) in the selected datasets.
Method 20% 30%
Random 20.80 19.83
MoSo 7.78 8.82
Moderate 0.30 0.31
Glister 21.21 21.95
Herding 35.00 30.56
Forgetting 23.00 21.76
GraNd 5.00 5.14
EL2N 22.00 21.80
SSP 21.70 20.21
CG-Score 45.09 39.69
Ours 0.24 0.25
@@
Figure 6: Effectiveness vs. efficiency on C-100. Results are reported with R-50 under 30% selection ratio on a 4-2080TI-GPU server.
@@
Training Efficiency Comparisons  To evaluate the selection efficiency of various methods, we present an analysis of the balance between effectiveness and efficiency. As shown in Figure 7, our method presents the best performance with desirable efficiency. Herding, EL2N, and GraNd obtain the lowest selection costs because they rely on predefined metrics or select samples in very early training. Our method is slightly slower than them but exhibits higher accuracy. Compared with the optimization-based approaches, like MoSo and Glister, our method enjoys both lower costs and better performance. The results verify the effectiveness of our method in balancing selection efficiency and accuracy.
@@ 4.3 Comparison of Robustness
As can be seen in Table 4.2, our approach exhibits superior robustness to noisy labels, significantly outperforming other baselines by a large margin. Specifically, our approach yields improvements of over 10.12% on CIFAR-100 and 4.41% on Tiny-ImageNet compared to previous leading methods. Additionally, Table 4.2 shows the distribution of noisy data within the selected datasets, where our method selects only 0.24% noisy samples, considerably fewer than other baselines. This reduction in noise underscores the potential of our method to improve overall data quality.
We argue that the robustness of our approach can be attributed to $\bm{S}_{A}$ in Eq. 1, which assesses the semantic alignment between image content and its labels. In the presence of label noise, this alignment is disrupted, resulting in a lower SAS, which in turn reduces the likelihood of such samples being selected during optimization. In contrast, most baseline methods rely solely on image features for selection, which may result in performance degradation when faced with noisy labels. In some cases, the performance of these methods is even worse than that of random selection. While Moderate also selects a relatively low proportion of noisy data, its performance is worse than ours. This discrepancy highlights the effectiveness of our method in making more strategic selections in noisy environments, thereby not only minimizing noises but also optimizing the selected datasets for improved generalization.
Robustness on Corrupted Images  We further evaluate the performance of our proposed method on real-world noise corruptions that are frequently encountered Singh et al. (2024). To simulate such corruptions, we employ the following five types of realistic noises Xia et al. (2023), including Gaussian noise, random occlusion, resolution, fog, and motion blur. The corruption rate is set to 5%, 10%, and 20%, respectively.
@@ 4.5 Ablation Study
Effect of Dataset Adaptation  To assess the effect of dataset adaptation, instead of using fine-tuned image and text adapters, we directly utilize the per-trained CLIP model to derive the scores $\bm{S}_{A}$ and $\bm{S}_{D}$. The experimental results, presented in Table 6 with a 90% selection ratio, show a significant decline in accuracy, with drops exceeding 2% on Tiny-ImageNet. Thus, dataset adaptation is essential for effectively transferring the model’s generalization ability to target datasets. This is particularly crucial for datasets that differ substantially from the pre-training datasets, such as CIFAR, where variations in image sizes and domain are distinct. Additionally, the adapters are implemented as MLP and fine-tuned with just 25 epochs, ensuring high efficiency.
@@
5 Conclusion

### v2

@@ Abstract
Large-scale datasets have been pivotal to the advancements of deep learning models in recent years, but training on such large datasets inevitably incurs substantial storage and computational overhead. Meanwhile, real-world datasets often contain redundant and noisy data, imposing a negative impact on training efficiency and model performance. Data selection has shown promise in identifying the most representative samples from the entire dataset, which aims to minimize the performance gap with reduced training costs. Existing works typically rely on single-modality information to assign importance scores for individual samples, which may lead to inaccurate assessments, especially when dealing with noisy or corrupted samples. To address this limitation, we propose a novel CLIP-powered data selection framework that leverages multimodal information for more robust and generalizable sample selection. Specifically, our framework consists of three key modules—dataset adaptation, sample scoring, and selection optimization—that together harness extensive pre-trained multimodal knowledge to comprehensively assess sample influence and optimize the selection results through multi-objective optimization. Extensive experiments demonstrate that our approach consistently outperforms existing state-of-the-art baselines on various benchmark datasets. Notably, our method effectively removes noisy or damaged samples from the dataset, enabling it to achieve even higher performance with less data. This indicates that it is not only a way to accelerate training but can also improve overall data quality. The implementation is available at https://github.com/Jackbrocp/clip-powered-data-selection.
@@ 1 Introduction
In this paper, we propose a CLIP-powered data selection framework that employs multimodal information for more robust and generalizable data selection, where the category text serves as a strong complement to the image modality and promotes overall performance. The framework consists of three modules, namely dataset adaptation, sampling scoring, and selection optimization module. First, the dataset adaptation module integrates image and text adapters to facilitate the transfer of pretraining knowledge to the target data. Subsequently, the sample scoring module calculates the Semantic Alignment Score (SAS) and Sample Diversity Score (SDS) based on the adapted multimodal features, which measure the image-text alignment and the variability of local patterns. Using these two scores together can select semantically representative samples while maintaining their inherent diversity. Further, in order to address the group effect, we introduce a selection optimization module to identify the ideal subsets w.r.t. the expected selection ratio through a multi-objective optimization strategy. By leveraging the multi-modal information and carefully designed supervision signals, our framework enables the selection of high-quality samples in a flexible and efficient manner.
Comprehensive evaluation across various benchmark datasets demonstrates that our approach effectively improves the performance of selected datasets, especially on large-scale datasets such as ImageNet-1k Deng et al. (2009). Moreover, the selected datasets exhibit superior cross-architecture generalization across ResNet-18/50 He et al. (2016), ViT Dosovitskiy et al. (2020), VGG-16 Simonyan & Zisserman (2014), etc. Notably, since most existing methods are not robust to more complex and realistic scenarios, we further validate the strong robustness of our method in more challenging scenes. For instance, our proposed method can achieve an 8.13% improvement in accuracy on CIFAR-100 and a 4.41% improvement on Tiny-ImageNet compared to the leading baselines. Meanwhile, superior performance is achieved with very high efficiency compared to other baselines.
@@ 2 Related Work
Date-efficient deep learning generally incorporates dynamic data pruning Qin et al. (2024); Raju et al. (2021), static data selection Xia et al. (2023); Tan et al. (2024); Yang et al. (2023b), data distillation Lei & Tao (2023); Du et al. (2023); Zhang et al. (2023); Cazenavette et al. (2022), and data condensation Liu et al. (2023); Yang et al. (2023a). Following the static data selection, we propose a method capable of identifying representative and diverse samples across various selection ratios. Data selection, or static dataset pruning, aims to identify and select the most representative samples from training datasets before training begins. Training on the selected subsets can achieve comparable performance to that obtained with the full dataset while reducing training and storage costs. Current data selection methods can be broadly divided into three categories: importance criteria Paul et al. (2021); Tan et al. (2024), dataset distribution-based methods Xia et al. (2023); Zheng et al. (2023), and optimization-based methods Nohyun et al. (2023); Yang et al. (2023c).
@@
Dataset distribution-based methods select samples by analyzing the geometric distribution of datasets. For instance, Herding Welling (2009) determines the sample importance according to samples’ distance to their corresponding class centers. The work Ramalingam et al. (2023) applies greedy k-center to select the coreset with good data coverage. D2 Maharana et al. (2023) calculates and updates the difficulty scores of each sample by incorporating the difficulty of its neighboring examples. Moderate-DS Xia et al. (2023) choose samples with closer distances to the median score, while CCS Zheng et al. (2023) balances the data distribution and the sample importance in selection.
Optimization-based methods select samples through various optimization techniques, such as temporal dual-depth scoring Zhang et al. (2024), gradient matching Mirzasoleiman et al. (2020b); Killamsetty et al. (2021a), scalable self-supervised pruning metric Sorscher et al. (2022), influence function Yang et al. (2023b); Pooladzandi et al. (2022), bi-level optimization Killamsetty et al. (2021b), facility location function Mirzasoleiman et al. (2020a); Yang et al. (2023d), and submodularity Iyer et al. (2021); Nohyun et al. (2023); Kothawade et al. (2022); Wei et al. (2015). In contrast to these methods that only rely on image information, we leverage multimodal messages for data selection, which incorporates the semantic alignment between image data and corresponding category information. This contributes to comprehensive assessments of sample effectiveness, particularly in complex scenarios where samples may be corrupted or wrongly labeled.
@@ 3.2 Sample Scoring
For selected datasets, the reduced data volume may limit the diversity of the selected data, which is important for training datasets Yang et al. (2024a). To address this, we introduce another diversity perspective to comprehensively assess the effect of training data. The Sample Diversity Score (SDS) is defined as the average distance between each sample and its $k$ neighbor samples of the same class:
@@ 3.3 Selection Optimization
where $\mathbb{I}$ is an indicator function, and $s_{r}$ denotes the expected selection ratio. The loss term guides the parameter $\bm{d}$ toward near-binary values, ensuring the count of ones aligns with the expected sample size. Since the selection is guided by adaptive optimization, the final selection ratio may deviate slightly from the target. To minimize this deviation, we constrain $\mathcal{L}_{s}$ with a threshold $\theta$, which in our work is set to $5\times 10^{-4}$, ensuring the actual selection ratio differs by less than ±$\pm 0.05\%$ from the expected value. We also provide a theoretical analysis of $\theta$ on the actual selection ratio gap in Appendix A. Finally, the overall loss function is formulated as:
@@
where $\alpha$ and $\beta$ are coefficients that adjust for numerical differences among the loss terms and can be set conveniently. The complete workflow is outlined in Algorithm 1 in Appendix B.
@@
The proposed method comprises three main components. 1) The dataset adaptation involves fine-tuning the image and text adapters. Since the adapters consist of simple linear layers, the number of parameters is small, and both the forward passes and backward passes are computationally efficient. 2) In the sample scoring process, the complexity of calculating the SAS and SDS is $O(N)$ and $O(K*f_{d})$, respectively, where $K$ is the number of categories and $f_{d}$ is the feature dimension, typically 512. The complexity of the KNN algorithm is $O(N_{k}*f_{d})$, where $N_{k}$ is the number of samples per class. Given that $K$ and $f_{d}$ are constants and $N_{k}$ is usually much smaller than $N$, the overall complexity of this process is approximately $O(N)$. 3) The selection optimization of $\bm{d}$ does not involve deep models and is a numerical optimization process. The complexity is proportional to the number of parameters, i.e., $O(|w|)=O(N)$.
@@
4 Experiment
4.1 Experimental Setup
Parameter settings. The parameters in our proposed method can be easily set. The coefficient $\alpha$ is proportional to the expected selection rate $s_{r}$, balancing the importance of dataset diversity, i.e., $\alpha$ can be set equivalent to $s_{r}$. The coefficient $\beta$ is set to 2 across datasets to adjust the numerical differences among loss items. For more details, please refer to Appendix C.
@@ 4.2 Comparison with the State-of-the-arts
Table 2: Experimental results (accuracy, %, mean $\pm$ std) on CIFAR-100 and Tiny-ImageNet with noisy labels. 20% of labels are disturbed. We also report the numerical analysis of the proportion (%) of noisy data in the selected CIFAR-100 datasets.
Method / Selection Ratio CIFAR-100 (label noise) Tiny-ImageNet (label noise) Noise ratios
20% 30% 20% 30% 20% 30%
Random 34.47$\pm$0.64 43.26$\pm$1.21 17.78$\pm$0.44 23.88$\pm$0.42 20.80 19.83
MoSo 31.01$\pm$0.67 43.73$\pm$0.14 21.55$\pm$0.37 27.80$\pm$0.16 7.78 8.82
Moderate-DS 40.25$\pm$0.12 48.53$\pm$1.60 19.64$\pm$0.40 24.96$\pm$0.30 0.30 0.31
Glister 28.51$\pm$1.46 43.16$\pm$1.31 21.61$\pm$0.19 25.45$\pm$0.23 21.21 21.95
Herding 42.29$\pm$1.75 50.52$\pm$3.38 18.98$\pm$0.44 24.23$\pm$0.29 35.00 30.56
Forgetting 36.53$\pm$1.11 45.78$\pm$1.04 13.20$\pm$0.38 21.79$\pm$0.43 23.00 21.76
GraNd 31.72$\pm$0.67 42.80$\pm$0.30 18.28$\pm$0.32 23.72$\pm$0.18 5.00 5.14
EL2N 29.82$\pm$1.19 33.62$\pm$2.35 13.93$\pm$0.69 18.57$\pm$0.31 22.00 21.80
Self-sup. prototypes 31.08$\pm$0.78 41.87$\pm$0.63 15.10$\pm$0.73 21.01$\pm$0.36 21.70 20.21
CG-Score 6.82$\pm$1.60 20.07$\pm$0.45 8.35$\pm$0.65 15.31$\pm$0.90 45.09 39.69
Ours 46.05$\pm$0.21 58.34$\pm$0.36 26.09$\pm$0.12 33.13$\pm$0.25 0.25 0.32
@@
Figure 6: Effectiveness vs. efficiency on C-100. Results are reported with R-50 under 30% selection ratio on a 4-2080TI GPU.
@@
Training Efficiency Comparisons  To evaluate the selection efficiency of various methods, we present an analysis of the balance between effectiveness and efficiency. As shown in Figure 7, our method presents the best performance with desirable efficiency. Herding, EL2N, and GraNd obtain the lowest selection costs because they rely on predefined metrics or select samples in very early training. Our method is slightly slower than theirs but exhibits higher accuracy. Compared with the optimization-based approaches, like MoSo and Glister, our method enjoys both lower costs and better performance. The results verify the effectiveness of our method in balancing selection efficiency and accuracy.
@@ 4.3 Comparison of Robustness
As can be seen in Table 2, our approach exhibits superior robustness to noisy labels, significantly outperforming other baselines by a large margin. Specifically, our approach yields improvements of over 10.12% on CIFAR-100 and 4.41% on Tiny-ImageNet compared to previous leading methods. Additionally, Table 2 shows the distribution of noisy data within the selected datasets, where our method selects only 0.24% noisy samples, considerably fewer than other baselines. This reduction in noise underscores the potential of our method to improve overall data quality.
We argue that the robustness of our approach can be attributed to $\bm{S}_{A}$ in Eq. 1, which assesses the semantic alignment between image content and its labels. In the presence of label noise, this alignment is disrupted, resulting in a lower SAS, which in turn reduces the likelihood of such samples being selected during optimization. In contrast, most baseline methods rely solely on image features for selection, which may result in performance degradation when faced with noisy labels. In some cases, the performance of these methods is even worse than that of random selection. While Moderate also selects a relatively low proportion of noisy data, its performance is worse than ours. This discrepancy highlights the effectiveness of our method in making more strategic selections in noisy environments, thereby not only minimizing noises but also optimizing the selected datasets.
Robustness on Corrupted Images  We further evaluate the performance of our proposed method on real-world noise corruptions that are frequently encountered Singh et al. (2024); Wei et al. (2024). To simulate such corruptions, we employ the following five types of realistic noises Xia et al. (2023), including Gaussian noise, random occlusion, resolution, fog, and motion blur. The corruption rate is set to 5%, 10%, and 20%, respectively.
@@ 4.5 Ablation Study
Effect of Dataset Adaptation  To assess the effect of dataset adaptation, instead of using fine-tuned image and text adapters, we directly utilize the per-trained CLIP model to derive the scores $\bm{S}_{A}$ and $\bm{S}_{D}$. The experimental results, presented in Table 5 with a 90% selection ratio, show a significant decline in accuracy, with drops exceeding 2% on Tiny-ImageNet. Thus, dataset adaptation is essential for effectively transferring the model’s generalization ability to target datasets. This is particularly crucial for datasets that differ substantially from the pre-training datasets, such as CIFAR, where variations in image sizes and domain are distinct.
Effect of Text Modality  Previous works Xia et al. (2023) use the average image features as the prototype and calculate the Euclidean distance between the embedded image and the corresponding prototype, which is used to filter noisy labels Wu et al. (2021). To evaluate the effect of our text modality, we also leverage this distance to assess the noisy samples and replace the text feature in Eq. 1. With a 20% noisy ratio, the accuracy drops from 46.05% to 16.39% with a 20% selection ratio and from 58.34% to 38.35% with a 30% selection ratio, validating the effectiveness of introducing textual modality. Due to the limited space, full results can be seen in Appendix L.
@@
5 Discussion and Conclusion
Limitation and future work. In this section, we discuss some potential limitations and future work for our method. 1). While the pretrained CLIP model contributes to selecting the most representative samples, the potential biases in the CLIP model may propagate to the selected dataset Alabdulmohsin et al. (2024). Future work may explore bias mitigation strategies by fine-tuning CLIP with bias-aware loss functions. 2). Our work primarily focuses on vision-based datasets for data selection, using textual modalities to guide sample selection, where both modalities are balanced. However, datasets with high modality imbalance may bring challenges since it is difficult to assess the modality alignment. Thus, future work could consider leveraging dynamic modality weight allocation mechanisms and augmentation strategies for modality-imbalanced dataset selection.